two cubes of identical volume but differing density are weighed. The first cube has a mass of 6.30 grams and the second cube has a mass of 3.44 grams. The density of the first cube is 1.55g/mL. What is the density of the second cube?
Answer #1

m1 = 6.3 g

m2 = 3.44 g

D1 = 1.55 g/ml

V1 = m1/D1 = 6.3/1.55 = 4.06451 ml

since V1 =V2 = 4.0645 ml

then

D2 = m2/v2 = 3.44 /4.064516129 = 0.8463 g/ml